OpenStudy (anonymous):

I WILL FAN AND MEDAL. Check my answer. A cylinder has a radius of 8 cm and a height of 14 cm. What is the volume of the cylinder? 896 pi cm3?

OpenStudy (calculusxy):

The formula for finding the volume of a cylinder is: \[V = B \times h\] \[B\] represents the area of the circle :\[\pi \times r^2\] So you really have: \[(\pi \times 8^2) \times 14\]

OpenStudy (anonymous):

so is it correct?

OpenStudy (calculusxy):

Did you multiply and see if it is correct? I find it helpful to catch the mistake myself and correct it.

OpenStudy (calculusxy):

What was your product?

OpenStudy (anonymous):

I got it thanks.

OpenStudy (calculusxy):

You're welcome :) Did you catch your mistake?